Part of a machine moves with a velocity given by the expression:7t^2+0.8tmm/s.
Calculate the displacement over the first 5 seconds of motion.

Answer by Alan3354(69443)

Part of a machine moves with a velocity given by the expression:7t^2+0.8tmm/s.
Calculate the displacement over the first 5 seconds of motion.
-----------
velocity is the 1st derivative of position, so integrate the expression
--> F(t) = 7t^3/3 + 0.4t^2 mm
F(5) = 7*125/3 + 0.4*25 = 875/3 + 10
F(5) = 905/3 mms
---------------------
No initial conditions were given, so the constant of integration was assumed to be zero.